OpenStudy (sarahc):

Part 1: Create and provide the dimensions for two similar figures of your choosing. Part 2: What is the similarity ratio of these figures along with the ratio of their surface area and volume? Part 3: Show your work, either using the actual volumes or using the formula, that the volume ratio is true.

OpenStudy (ash2326):

Take two similar cuboids, let the first one have the dimensions \(4\times 6\times 8\) Now the ratio between the two (you can choose any), let it be 1/2 Now can you tell me the dimensions of the second cuboid

OpenStudy (sarahc):

can it be anything?

OpenStudy (ash2326):

nope, we chose the ratio as 1:2. and we have the dimensions of the first cuboid. Now use it to find the dimensions

OpenStudy (sarahc):

2x3x4? idk...so lost.

OpenStudy (ash2326):

the ratio is 1:2 let the length of the second be L and we have the first cuboid's length as 4, so \[\frac 4 L=\frac 12\] Now you can find the length from this, similarly find other dimensions, ratio would be the same for all. Do you get this?

OpenStudy (sarahc):

okay so 8?

OpenStudy (ash2326):

yeah, correct. Now find the other two dimensions

OpenStudy (sarahc):

how? sorry :?

OpenStudy (ash2326):

the same way as we found length, use the first one's width, you have the ratio and find the second cuboid's width

OpenStudy (sarahc):

I have a feeling I'm making this seem harder than it really is..

OpenStudy (ash2326):

it's easy, i'll show you once more width of first one 6 let the width of second be w and we have the ratio as 1/2 so \[\frac 6 w=\frac 12\] now find w from here:)

OpenStudy (sarahc):

12..

OpenStudy (ash2326):

Yeah, now find the height:)

OpenStudy (sarahc):

16

OpenStudy (ash2326):

good :D Did you understand till here?

OpenStudy (sarahc):

okay so that would be the demotions of the other one?

OpenStudy (sarahc):

Lol I'm half asleep, sorry

OpenStudy (sarahc):

dimensions

OpenStudy (ash2326):

no worries, you should sleep now. I'll help when you are awake

OpenStudy (ash2326):

yeah the dimensions of the other cuboid, the two cuboids are similar with a similarity ratio of 1/2

OpenStudy (sarahc):

Okay :) the ratio of their surface area and volume? :o

OpenStudy (ash2326):

@sarahc I gotta bounce now, Sorry my friends @lgbasallote @apoorvk will help you out

OpenStudy (sarahc):

okay...thanks for your help !